On June 1, 2016, Zinba Company retired a piece of equipment that was originally purchased for $250,000. At December 31, 2015, the equipment had a book value of $75,000 and was being depreciated on a straight-line basis at $25,000 per year. Zinba sold the equipment for $60,000.


Prepare the journal entries necessary to record the disposal of the machine. You must SHOW YOUR WORK!!!
Answer #1

Journal entry on disposal of machine

Debit Credit
Cash   $60,000   
Accumulated Depreciation $175,000
Loss on disposal of Equipment $15,000
Equipment $250,000

Accumulated Depreciation = 250,000-75000 = 175,000

Loss on disposal of Equipment = 250,000- 175,0000-60,000 = 15,000